# Graph Masked Autoencoder for Spatio-Temporal Graph Learning

### v1

@@ Abstract
Effective spatio-temporal prediction frameworks play a crucial role in urban sensing applications, including traffic analysis, human mobility behavior modeling, and citywide crime prediction. However, the presence of data noise and label sparsity in spatio-temporal data presents significant challenges for existing neural network models in learning effective and robust region representations. To address these challenges, we propose a novel spatio-temporal graph masked autoencoder paradigm that explores generative self-supervised learning for effective spatio-temporal data augmentation. Our proposed framework introduces a spatial-temporal heterogeneous graph neural encoder that captures region-wise dependencies from heterogeneous data sources, enabling the modeling of diverse spatial dependencies. In our spatio-temporal self-supervised learning paradigm, we incorporate a masked autoencoding mechanism on node representations and structures. This mechanism automatically distills heterogeneous spatio-temporal dependencies across regions over time, enhancing the learning process of dynamic region-wise spatial correlations. To validate the effectiveness of our STGMAE framework, we conduct extensive experiments on various spatio-temporal mining tasks. We compare our approach against state-of-the-art baselines. The results of these evaluations demonstrate the superiority of our proposed framework in terms of performance and its ability to address the challenges of spatial and temporal data noise and sparsity in practical urban sensing scenarios.
Index Terms:
Graph Autoencoder, Spatio-Temporal Graph Learning
I Introduction
@@
As remote sensing technologies and large-scale computing infrastructures continue to provide an abundance of spatio-temporal data, there is an increasing demand for robust spatio-temporal prediction frameworks in urban sensing applications. These applications include traffic analysis, such as the works by STGODE,  [1, 2, 3, 4] as well as modeling human mobility behavior, as explored by studies  [5, 6, 7]. Additionally, citywide crime prediction, as studied by Chen and Diao [8, 9], is another important area of focus. A crucial aspect of developing effective spatio-temporal prediction frameworks lies in the accurate capture of spatial and temporal correlations between different geographical locations and time intervals. This ability ensures comprehensive insights into dynamic urban phenomena and supports informed decision-making processes.
@@
Researchers have introduced spatio-temporal graph neural networks (STGNNs) as a promising learning paradigm. Notable works in this area include [10, 11, 12]. These networks are specifically designed to handle data with graph structures and have garnered significant attention due to their ability to capture multi-hop spatial and temporal dependencies. This is accomplished through the iterative propagation of messages among nodes, leveraging their knowledge of both spatial and temporal information. By employing this recursive message passing mechanism, STGNNs demonstrate exceptional proficiency in modeling intricate dependencies in spatio-temporal data. Their effectiveness lies in their capability to integrate and exploit the inherent structure and temporal dynamics of the data, enabling more accurate predictions and insightful analysis of complex spatio-temporal phenomena in various real-world applications.
@@
However, existing works in STGNNs have not adequately addressed the challenges of noise and sparsity in spatio-temporal data, which hinders the accurate modeling of dependencies. For instance, traffic data collected from sensors or GPS devices may contain outliers or missing values, leading to less relevant region connections and noisy signals. These noisy signals can negatively impact the performance of spatio-temporal prediction tasks, as they introduce irrelevant information and hinder the accurate modeling of dependencies. Besides, most existing solutions approach spatio-temporal prediction tasks in a supervised manner, relying on a sufficient amount of labeled data for training. However, spatio-temporal data is often sparsely labeled, especially for rare urban anomalies or events. This sparsity presents a significant challenge in building accurate prediction models. Thus, it becomes crucial to develop learning models that can effectively handle sparse labeled data and learn from limited samples for achieving accurate spatio-temporal predictions in real-world urban scenarios.
@@
In this paper, we propose Spatio-Temporal Graph Masked Autoencoder (STGMAE) to address the challenges of noise and sparsity in spatio-temporal data. Our framework leverages multi-view spatio-temporal data and introduces a spatio-temporal heterogeneous graph neural encoder to capture region-wise dependencies. By considering both intra-view and inter-view data correlations, our model effectively models spatio-temporal patterns. Moreover, we propose a self-supervised learning paradigm that utilizes the mask mechanism to distill heterogeneous dependencies across observed regions in urban space. Specifically, we adopt the masked autoencoding mechanism on the node representations and graph structures to identify and emphasize the most relevant connections and features for prediction tasks. This enhances the learning process of implicit cross-region relations and improves the robustness of the model against noisy signals.
@@
At the core of our approach is a generative spatio-temporal graph auto-encoder. This model learns to reconstruct both node features and the graph structure by utilizing the mask mechanism. By selectively masking certain regions in the input data during training, our model focuses on capturing the dependencies between regions and effectively reconstructing the missing or corrupted features and connections. This capability effectively overcomes the data sparsity by providing accurate self-supervisory signals to prioritize the most relevant information for spatio-temporal relation learning.
@@
Our STGMAE framework has been extensively evaluated in diverse spatio-temporal mining tasks and compared against multiple region representation and GNN baselines. The results of these evaluations unequivocally demonstrate the effectiveness and superiority of our proposed method in addressing the challenges associated with data noise. We highlight the key contributions as follows:
@@
  • We place a strong emphasis on addressing the challenges of data noise in spatio-temporal prediction tasks. To achieve this, we introduce the superiority of generative contrastive learning in region representation, incorporating the mask mechanism to enhance the robustness and generalization ability of our spatio-temporal heterogeneous graph neural architecture.
@@
  • We propose a spatio-temporal heterogeneous graph neural network that captures both intra-view and inter-view region dependencies while modeling various spatio-temporal factors. To address the data sparsity issue, we perform generative spatio-temporal self-supervised learning over the region dependency graph. Our framework is built upon the masked autoencoding mechanism via masking node features and the graph structure on the heterogeneous spatio-temporal graph.
@@
  • We rigorously validate our STGMAE across a range of spatio-temporal mining tasks, conducting comprehensive comparisons with state-of-the-art baselines. Our method consistently achieves the best performance across all tasks, including crime prediction and traffic prediction.
II Related Work
Self-supervised Learning for Spatio-Temporal Graphs. In the field of self-supervised learning (SSL) for spatio-temporal graphs, research has focused on capturing spatial and temporal dependencies within the graph structure. Recent studies [13, 14, 15, 16] propose innovative approaches in this area. Liu et al. [13] address data sparsity in POI recommendation by leveraging supplementary information through SSL. Ji et al. [14] enhance traffic prediction with SSL, capturing spatial and temporal heterogeneity. Zhang et al. [15] improve region representations using SSL, capturing underlying patterns and dynamics. Zhang et al. [16] combine SSL with adversarial training, enhancing graph representation and downstream task performance. These studies demonstrate the effectiveness of SSL in capturing spatio-temporal dependencies and improving representation learning for various tasks such as POI recommendation, traffic prediction, and region representations.
Graph Mask Autoencoder. Graph autoencoder methods are generative self-supervised learning approaches that construct input graphs directly, eliminating the need for data augmentation. Kipf and Welling [17] introduce Graph Autoencoder (GAE) and Variational Graph Autoencoder (VGAE) using Graph Convolutional Networks (GCN) for topology and node content reconstruction. ARVGA [18] proposes an adversarial graph embedding framework, enhancing robustness and generalization. MGAE [19] and GALA [20] combine structure and node reconstruction for a comprehensive graph representation. GraphMAE [21] focuses on node reconstruction, achieving promising results in node classification. Overall, graph autoencoders offer a powerful framework for self-supervised learning on graphs, excelling in tasks like graph generation and node classification.
Region Representation Learning. Research on the region representation learning problem includes the following studies [22, 23, 24, 25, 26, 27, 15]. In particular, Fu et al. [26] suggest to incorporate both intra-region information (such as POI distance inside a region) and inter-region information (such as the similarity of POI distributions between two regions) to increase the quality of region representations. The concept of putting forward in [26] is expanded upon by Zhang et al. [25] by using a collective adversarial training.
In a recent study conducted by Zhang et al. [24], a multi-view joint learning model is developed to learn region representations. To capture region correlations from different perspectives, such as human mobility and region properties, the study incorporates a graph attention technique for each view. However, this approach focuses solely on modeling region correlations and does not consider the inclusion of point-of-interest (POI) data, which is essential for capturing region functionality. Similarly, Wu et al. [27] propose a method for learning area representations by extracting traffic patterns. However, their approach also excludes POI data and relies solely on mobility data. The efficacy of these techniques heavily relies on the generation of high-quality region graphs. Furthermore, they may struggle to learn accurate region representations when faced with noisy and skewed spatio-temporal data. In contrast, a recent study by Zhang et al. [15] proposes the adoption of a contrastive self-supervised learning paradigm on a multi-view heterogeneous spatio-temporal graph. This approach utilizes a self-contrastive learning method to obtain high-quality region vectors. By contrast, our paper employs a generative self-supervised learning method to model region correlations by masking links and nodes on the heterogeneous spatio-temporal graph.
III Methodology
@@
Figure 1: Architecture of our STGMAE region representation model with graph masked autoencoder.
@@
The framework of STGMAE is depicted in Figure 1, which addresses the task of reconstructing links and features of masked nodes using a mask mechanism based on generative self-supervised learning on the heterogeneous spatio-temporal graph. The STGMAE framework consists of an encoder and a decoder. The encoder utilizes GCN [28] to capture the underlying graph structure and extract meaningful representations from the heterogeneous spatio-temporal graph $\mathcal{G}$. The decoder, also employing GCN, is responsible for reconstructing the masked links and features of nodes in the graph. The STGMAE framework operates on the heterogeneous spatio-temporal graph $\mathcal{G}$, which incorporates various types of information, such as spatial, temporal, and attribute data. By utilizing the mask mechanism, STGMAE learns to reconstruct the missing links and features of masked nodes, thereby capturing the correlations and dependencies. And the specific algorithm is shown in Algorithm 1.
III-A Preliminaries
@@
In our methodology, we employ a partitioning approach to divide the geographic area of a city (such as New York) into $I$ spatio-regions. Each spatio-region is identified by an index, denoted as $i$ (e.g., $r_{i}$). To capture a wide range of urban contextual information, we propose a region embedding framework that integrates multiple diverse data sources associated with the metropolitan space. These data sources serve as inputs to our STGMAE, and enriched representation of each spatio-region.
@@
Human Mobility Trajectories. We create a matrix $\mathbf{M}\in\mathbb{R}^{I\times I}$ utilizing human trajectory to accurately represent the dynamic urban flows between citywide zones, where $T$ is the number of time slots. Individual human mobility traces are saved in this collection as $(r_{s},r_{d},t_{s},t_{d})$, where $r_{s}$ and $r_{d}$ stand in for source and destination regions, respectively, and $t_{s}$ and $t_{d}$ stand for timestamp data.
@@
Spatio Correlations. We create a spatio matrix $\mathcal{D}\in\mathbb{R}^{I\times T}$. It is composed of region pairs like $(r_{s},r_{d},d^{\prime})$ whose geographical distances are less than a threshold $\epsilon$, where $d^{\prime}$ denotes the distance from the source node $r_{s}$ to the destination node $r_{d}$.
@@
Problem Statement: Given the spatio-temporal graph, denoted as $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ and $\mathcal{E}$ are node set and edge set respectively. $\mathcal{G}$ is built upon the POI matrix $\mathbf{P}$, the mobility matrix $\mathbf{M}$ and the spatio matrix $\mathcal{D}$. The task of this work is to learn a function $f$, which maps each node of $\mathcal{G}$ to the embedding space and reconstruct both the structure $\mathbf{A}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|}$ and features $\mathbf{X}\in\mathbb{R}^{N\times D}$ of the input spatio-temporal graph $\mathcal{G}$. Meanwhile, the embedding space can effectively preserve the geographical and temporal connections between various locations and time slots, which is advantageous for a variety of urban sensing applications (e.g.traffic forecasting and crime prediction).
III-B Embedding Layer
@@
To incorporate the Point-of-Interest (POI) context into region embeddings, we have developed a POI context embedding layer that reflects the functional information of each region in the latent representation space. Inspired by the POI token embedding technique using Skip-gram [29], we input the region-specific POI vector into the Skip-gram model [30] for POI context embedding. Subsequently, we concatenate the region-specific POI embeddings to generate POI-aware representations: $\bar{\textbf{E}}=\text{MLP}(\text{Skip-gram}(\mathcal{P}))$. Here, we employ a Multi-Layer Perceptron (MLP) for the embedding projection. The embedding table $\textbf{E}\in\mathbb{R}^{I\times d}$ represents the embeddings for all regions, where $d$ is the hidden dimensionality. Each region’s embedding $\textbf{e}_{i}$ captures the region-specific POI contextual information in the latent semantic space. To incorporate region interactions within our POI embedding layer, we introduce a region-wise self-attention mechanism given by:
@@
$\textbf{e}_{i}=\mathop{\Bigm{|}\Bigm{|}}_{h=1}^{H}\sum_{j=1}^{I}\alpha_{i,j}^{h}\cdot\textbf{V}^{h}\textbf{e}_{j};\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \alpha_{i,j}^{h}=\delta\left(\frac{(\textbf{Q}^{h}\textbf{e}_{i})^{\top}(\textbf{K}^{h}\textbf{e}_{j})}{\sqrt{d/H}}\right)$ (1)
@@
III-C Heterogeneous Spatio-Temporal GNNs
@@
Our methods begins with designing a spatio-temporal graph $\mathcal{G}$, which is composed of three matrixes, namely region POI matrix $\mathbf{P}$, human trajectory matrix $\mathbf{M}$ and spatio distance matrix $\mathcal{D}$. Thus the relations of the graph $\mathcal{G}$ are classified into five types, namely relations $\mathcal{R}_{p}$ based on POI similarities of regions, relations $\mathcal{R}_{m}$ based on urban flow between regions, relations $\mathcal{R}_{d}$ based on distances of regions, relations $\mathcal{R}_{pm}$ between the POI subgraph and the mobility subgraph, and relations $\mathcal{R}_{md}$ between the mobility subgraph and spatio distance graph. And the relation set $\mathcal{R}$ is shown as $\mathcal{R}=(\mathcal{R}_{p},\mathcal{R}_{m},\mathcal{R}_{d},\mathcal{R}_{pm},\mathcal{R}_{md})$.Thus, relation-aware message passing paradigm is represented as:
@@
$\textbf{h}_{i}^{(l+1)}=\sigma(\sum_{\gamma\in\mathcal{R}}\sum_{j\in\mathcal{N}^{\gamma}_{i}}\alpha_{i,\gamma}\textbf{W}_{\gamma}^{(l)}\textbf{h}_{j}^{(l)});\leavevmode\nobreak\ \leavevmode\nobreak\ \alpha_{i,\gamma}=\frac{1}{|\mathcal{N}^{\gamma}_{i}|}$ (2)
@@
where $\mathcal{N}^{\gamma}_{i}$ denotes neighbour set of region $r_{i}$ via relation $\gamma$ ($\gamma\in\Gamma$). $\mathcal{R}$ is the relation set. $\textbf{h}_{i}^{(l+1)},\textbf{h}_{j}^{(l)}\in\mathbb{R}^{d}$ represents the hidden embedding vectors of the $i$-th region vertex in the $(l+1)$-th graph layer and the $j$-th region vertex in the $l$-th graph neural layer, respectively. $\sigma(\cdot)$ denotes the ReLU activation function. $\alpha_{i,\gamma}\in\mathbb{R}$ represents the normalization weight of region vertex pair $(r_{i},r_{j})$, which is calculated using the degrees of the vertices. $\textbf{W}_{\gamma}^{(l)}\in\mathbb{R}^{d\times d}$ presents the learning weights for the $(l)-th$ iteration. STGMAE combines the multi-order embeddings to fully utilize the data obtained from the multi-hop neighborhood. The cross-view message passing mechanism is shown in the following formula:
$\textbf{H}=\sum_{l=0}^{L}\textbf{H}^{(l)};\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \textbf{H}^{(l+1)}=\sigma(\textbf{D}^{-\frac{1}{2}}\textbf{A}\textbf{D}^{-\frac{1}{2}}\textbf{H}^{(l)}\textbf{W}_{\gamma}^{(l+1)\top})$ (3)
@@
where $\textbf{H}\in\mathbb{R}^{|\mathcal{V}|\times d}$ is the embedding matrix, whose rows are regional embedding vectors $\textbf{h}_{i}$. $L$ is the number of graph iterations. $\textbf{A}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|}$ denotes the adjacent matrix with self-loop. The diagonal degree matrix is denoted as $\textbf{D}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|}$.
III-D The Layout of Spatial-Temporal Graph Mask Autoencoder (STGMAE)

### v2

# HGAurban: Heterogeneous Graph Autoencoding for Urban Spatial-Temporal Learning

@@ Abstract.
Spatial-temporal graph representations play a crucial role in urban sensing applications, including traffic analysis, human mobility behavior modeling, and citywide crime prediction. However, a key challenge lies in the noisy and sparse nature of spatial-temporal data, which limits existing neural networks’ ability to learn meaningful region representations in the spatial-temporal graph. To overcome these limitations, we propose HGAurban, a novel heterogeneous spatial-temporal graph masked autoencoder that leverages generative self-supervised learning for robust urban data representation. Our framework introduces a spatial-temporal heterogeneous graph encoder that extracts region-wise dependencies from multi-source data, enabling comprehensive modeling of diverse spatial relationships. Within our self-supervised learning paradigm, we implement a masked autoencoder that jointly processes node features and graph structure. This approach automatically learns heterogeneous spatial-temporal patterns across regions, significantly improving the representation of dynamic temporal correlations. Comprehensive experiments across multiple spatiotemporal mining tasks demonstrate that our framework outperforms state-of-the-art methods and robustly handles real-world urban data challenges, including noise and sparsity in both spatial and temporal dimensions.
1. Introduction
@@
As remote sensing technologies and large-scale computing infrastructures continue to provide an abundance of spatial-temporal data, there is an increasing demand for robust spatial-temporal prediction frameworks in urban sensing applications. These applications include traffic analysis (Fang et al., 2021; Wang et al., 2020; Li and Zhu, 2021; Wang et al., 2021; Zhang et al., 2025b, d; Wang et al., 2019; Lv et al., 2018; Zhang et al., 2025a), as well as modeling human mobility behavior (Pentland and Liu, 1999; Fuchs et al., 2023; Zhang et al., 2024a, 2023c). Additionally, citywide crime prediction (Chen et al., 2021; Diao et al., 2019; Long et al., 2023; Zhang et al., 2025c, 2024b) is another important area of focus. A crucial aspect of developing effective spatial-temporal prediction frameworks lies in the accurate capture of spatial and temporal correlations between different geographical locations and time intervals. This ability ensures comprehensive insights into dynamic urban phenomena.
@@
Researchers have introduced spatial-temporal graph neural networks (STGNNs) as a promising learning paradigm (Sahili and Awad, 2023; Jin et al., 2023; Ta et al., 2022). These networks are specifically designed to handle data with graph structures and have garnered significant attention due to their ability to capture multi-hop spatial and temporal dependencies. This is accomplished through the iterative propagation of messages among nodes, leveraging their knowledge of both spatial and temporal information. By employing this recursive message passing mechanism, STGNNs demonstrate exceptional proficiency in modeling intricate dependencies in spatial-temporal data. Their effectiveness lies in their capability to integrate and exploit the inherent structure and temporal dynamics of the data, enabling more accurate predictions and insightful analysis of various urban applications.
@@
However, existing works in STGNNs have not adequately addressed the challenges of noise and sparsity in spatial-temporal data, which hinders the accurate modeling of dependencies. For instance, traffic data collected from sensors or GPS devices may contain outliers or missing values, leading to less relevant region connections and noisy signals. These noisy signals can negatively impact the performance of spatial-temporal prediction tasks, as they introduce irrelevant information and hinder the accurate modeling of dependencies. Besides, most existing solutions approach spatial-temporal prediction tasks in a supervised manner, relying on a sufficient amount of labeled data for training. However, spatial-temporal data is often sparsely labeled, especially for rare urban anomalies or events. This sparsity presents a significant challenge in building accurate prediction models. Thus, it becomes crucial to develop learning models that can effectively handle sparse labeled data and learn from limited samples for achieving accurate spatial-temporal predictions in real-world urban scenarios.
@@
In light of these challenges, we propose heterogeneous graph autoencoding for urban spatial-temporal learning (HGAurban). Our framework leverages multi-view spatial-temporal data and introduces a spatial-temporal heterogeneous graph neural encoder to capture region-wise dependencies. By considering both intra-view and inter-view data correlations, our model effectively models spatial-temporal patterns. Moreover, we propose a self-supervised learning paradigm that utilizes the mask mechanism to distill heterogeneous dependencies across observed regions in urban space. Specifically, we adopt the masked autoencoding mechanism on the node representations and graph structures to identify and emphasize the most relevant connections for prediction tasks. This enhances the learning process of implicit cross-region relations and improves the robustness of the model against noisy signals.
@@
At the core of our approach is a generative spatial-temporal graph auto-encoder. This model learns to reconstruct both node features and the graph structure by utilizing the mask mechanism. By selectively masking certain regions in the input data during training, our model focuses on capturing the dependencies between regions and effectively reconstructing the missing or corrupted features and connections. This capability effectively overcomes the data sparsity by providing accurate self-supervisory signals to prioritize the most relevant information.
@@
Our HGAurban framework has been extensively evaluated in diverse spatial-temporal mining tasks and compared against multiple region representation and GNN baselines. The results demonstrate the effectiveness and superiority of our proposed method in addressing the challenges associated with data noise and sparsity. We highlight the key contributions as follows:
@@
  • We place a strong emphasis on addressing the challenges of data noise in spatial-temporal prediction tasks. To achieve this, we introduce the superiority of self-supervised learning in region representation, incorporating the mask mechanism to enhance the robustness and generalization ability of our spatial-temporal heterogeneous graph neural architecture.
@@
  • We propose a spatiotemporal heterogeneous graph neural network that captures both intra-view and inter-view regional dependencies while modeling diverse spatiotemporal factors. To address data sparsity, we introduce a masked autoencoding mechanism that masks both node features and graph structures within the heterogeneous spatiotemporal graph, enabling the model to effectively explore inter-correlations among regions.
@@
  • We rigorously validate our HGAurban across a range of spatial-temporal mining tasks, conducting comprehensive comparisons with state-of-the-art baselines. Our method consistently achieves the best performance across three diverse tasks. The code is publicly available at https://github.com/lizzyhku/HGAurban.
2. Methodology
@@
Figure 1. Architecture of our HGAurban region representation model with graph masked autoencoder.
@@
The framework of HGAurban is depicted in Figure 1, which addresses the task of reconstructing links and features of masked nodes using a mask mechanism based on generative self-supervised learning on the heterogeneous spatial-temporal graph. The HGAurban framework operates on the heterogeneous spatial-temporal graph $\mathcal{G}$, which incorporates various types of information, such as spatial, temporal, and attribute data. Our HGAurban employs an encoder-decoder architecture. The encoder captures the underlying structure and derives meaningful representations from the heterogeneous spatiotemporal graph $\mathcal{G}$, while the decoder reconstructs the masked node features and graph links. By utilizing the mask mechanism, HGAurban learns to reconstruct the missing links and features of masked nodes, thereby capturing the correlations and dependencies. And the specific algorithm of HGAurban is shown in Algorithm 1.
2.1. Preliminaries
@@
In our methodology, we employ a partitioning approach to divide the geographic area of a city (such as New York) into $I$ spatial-regions. Each spatial-region is identified by an index, denoted as $i$ (e.g., $r_{i}$). To capture a wide range of urban contextual information, we propose a region embedding framework that integrates multiple diverse data sources associated with the metropolitan space. These data sources are incorporated into HGAurban to generate enriched spatial-region representations, which are shown as follows.
@@
Human Mobility Trajectories. We create a matrix $\mathbf{M}\in\mathbb{R}^{I\times T}$ utilizing human trajectory to accurately represent the dynamic urban flows between citywide zones, where $T$ is the number of time slots. Individual human mobility traces are saved in this collection as $(r_{s},r_{d},t_{s},t_{d})$, where $r_{s}$ and $r_{d}$ denote source and destination regions, respectively, and $t_{s}$ and $t_{d}$ stand for timestamp data.
@@
Spatial Correlations. We create a spatial matrix $\mathcal{D}\in\mathbb{R}^{I\times I}$. It is composed of region pairs like $(r_{s},r_{d},d^{\prime})$ whose geographical distances are less than a threshold $\epsilon$, where $d^{\prime}$ denotes the distance from the source node $r_{s}$ to the destination node $r_{d}$.
@@
Problem Statement: Given the spatial-temporal graph, denoted as $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ and $\mathcal{E}$ are node set and edge set respectively. $\mathcal{G}$ is built upon the POI matrix $\mathcal{P}$, the mobility matrix $\mathbf{M}$ and the spatial matrix $\mathcal{D}$. The task of this work is to learn a function $f$, which maps each node of $\mathcal{G}$ to the embedding space and reconstruct both the structure $\mathbf{A}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|}$ and features $\mathbf{X}\in\mathbb{R}^{|\mathcal{V}|\times D}$ of the input spatial-temporal graph $\mathcal{G}$. Meanwhile, the embedding space can effectively preserve the geographical and temporal connections between various locations and time slots, which is advantageous for a variety of urban sensing applications (e.g.traffic forecasting and crime prediction).
2.2. Embedding Layer
@@
To incorporate the Point-of-Interest (POI) context into region embeddings, we have developed a POI context embedding layer that reflects the functional information of each region in the latent representation space. Inspired by the POI token embedding technique using Skip-gram (Rahmani et al., 2019), we input the region-specific POI vector into the Skip-gram model (Cheng et al., 2006) for POI context embedding. Subsequently, we concatenate the region-specific POI embeddings to generate POI-aware representations: $\textbf{E}=\text{MLP}(\text{Skip-gram}(\mathcal{P}))$.
Here, we employ a Multi-Layer Perceptron (MLP) for the embedding projection. The embedding table $\textbf{E}\in\mathbb{R}^{I\times d}$ represents the embeddings for all regions, where $d$ is the hidden dimensionality. Each region’s embedding $\textbf{e}_{i}$ captures the region-specific POI contextual information in the latent semantic space. To incorporate region interactions within our POI embedding layer, we introduce a region-wise self-attention mechanism given by:
@@
(1) $\textbf{e}_{i}=\mathop{\Bigm{|}\Bigm{|}}_{h=1}^{H}\sum_{j=1}^{I}\alpha_{i,j}^{h}\cdot\textbf{V}^{h}\textbf{e}_{j};\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \alpha_{i,j}^{h}=\delta\left(\frac{(\textbf{Q}^{h}\textbf{e}_{i})^{\top}(\textbf{K}^{h}\textbf{e}_{j})}{\sqrt{d/H}}\right)$
@@
2.3. Heterogeneous Spatial-Temporal GNNs
@@
We model our data as a spatial-temporal graph $\mathcal{G}$, which is composed of three matrices, namely region POI matrix $\mathbf{P}$, human trajectory matrix $\mathbf{M}$ and spatial distance matrix $\mathcal{D}$. Thus the relations of the nodes in graph $\mathcal{G}$ are diverse and can be classified into five types, namely relations $\mathcal{R}_{p}$ based on POI similarities of regions, relations $\mathcal{R}_{m}$ based on urban flow between regions, relations $\mathcal{R}_{d}$ based on distances of regions, relations $\mathcal{R}_{pm}$ between the POI subgraph and the mobility subgraph, and relations $\mathcal{R}_{md}$ between the mobility subgraph and spatial distance graph. And the relation set $\mathcal{R}$ is shown as $\mathcal{R}=\{\mathcal{R}_{p},\mathcal{R}_{m},\mathcal{R}_{d},\mathcal{R}_{pm},\mathcal{R}_{md}\}$. To handle the heterogeneous relationship, we design the relation-aware message passing paradigm as follows:
@@
(2) $\textbf{h}_{i}^{(l+1)}=\sigma(\sum_{\gamma\in\mathcal{R}}\sum_{j\in\mathcal{N}^{\gamma}_{i}}\alpha_{i,\gamma}\textbf{W}_{\gamma}^{(l)}\textbf{h}_{j}^{(l)});\leavevmode\nobreak\ \leavevmode\nobreak\ \alpha_{i,\gamma}=\frac{1}{|\mathcal{N}^{\gamma}_{i}|}$
@@
where $\mathcal{N}^{\gamma}_{i}$ denotes neighbour set of region $r_{i}$ via relation $\gamma$. $\mathcal{R}$ is the relation set. $\textbf{h}_{i}^{(l+1)},\textbf{h}_{j}^{(l)}\in\mathbb{R}^{d}$ represents the hidden embedding vectors of the $i$-th region vertex in the $(l+1)$-th graph layer and the $j$-th region vertex in the $l$-th graph neural layer, respectively. $\sigma(\cdot)$ denotes the ReLU activation function. $\alpha_{i,\gamma}\in\mathbb{R}$ represents the normalization weight of region vertex pair $(r_{i},r_{j})$, which is calculated using the degrees of the vertices. $\textbf{W}_{\gamma}^{(l)}\in\mathbb{R}^{d\times d}$ presents the learning weights for the lthl-th layer. HGAurban combines the multi-order embeddings to fully utilize the data obtained from the multi-hop neighborhood. The cross-view message passing mechanism is shown in the following formula:
(3) $\textbf{H}=\sum_{l=0}^{L}\textbf{H}^{(l)};\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \textbf{H}^{(l+1)}=\sigma(\textbf{D}^{-\frac{1}{2}}\textbf{A}\textbf{D}^{-\frac{1}{2}}\textbf{H}^{(l)}\textbf{W}_{\gamma}^{(l+1)\top})$
@@
where $\textbf{H}\in\mathbb{R}^{|\mathcal{V}|\times d}$ is the embedding matrix, whose rows are regional embedding vectors $\textbf{h}_{i}$. $L$ is the number of the model layers. $\textbf{A}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|}$ denotes the adjacent matrix with self-loop. The diagonal degree matrix is denoted as $\textbf{D}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|}$.
2.4. The Layout of Spatial-Temporal Graph Mask Autoencoder